OpenStudy (anonymous):

Step by step. To solve for x: 2x^2+6x+7=0

OpenStudy (anonymous):

You need to use the Quadratic Formula

OpenStudy (anonymous):

\[\frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] where a=2 b=6 c = 7

OpenStudy (kirbykirby):

\[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}=\frac{-6 \pm \sqrt{6^2-4(2)(7)}}{2(2)}\]

OpenStudy (kirbykirby):

You won't have an real solutions though since the discriminant is negative

OpenStudy (anonymous):

\[\frac{- 6 \pm \sqrt{6^2 - 4(2)(7)}}{4}\]\[\frac{-6 \pm \sqrt{-20}}{4}\]\[\frac{-6 \pm 2i \sqrt{5}}{4}\]\[\frac{-3 \pm \sqrt{5}i}{2}\]

OpenStudy (anonymous):

oh, now I understand! Thank you both you've been a big help, I really appreciate it.